# Composing Codensity Bisimulations

### v1

@@ 5 Codensity Liftings of structure Functors
We use the generalized codensity lifting to lift the structure functor. Let $(\Omega,\mathbf{\Omega}):Id_{A}\to(p:\mathbb{E}\to\mathbb{B})$ be a ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values. Its $N$-fold tupling $(\langle\Omega\rangle_{N},\langle\mathbf{\Omega}\rangle_{N}):Id_{A}\to(p^{N}:\mathbb{E}^{N}\to\mathbb{B}^{N})$ is again a ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values; here $\langle\Omega\rangle_{N}$ denotes the tupling $\langle\Omega,\cdots,\Omega\rangle:A\to\mathbb{B}^{N}$ of $N$-fold copies of $\Omega$ (and the same for $\langle\mathbf{\Omega}\rangle_{N}$). We call the codensity lifting of a structure functor $T:\mathbb{B}^{N}\to\mathbb{B}$ with $\sigma\colon T\circ\langle\Omega\rangle_{N}\to\Omega$ along $p^{N},p$ the $N$-codensity lifting. Concretely,
@@ 5.1 Product Functors by Codensity Liftings
One of the most fundamental operators for composing processes and state transition systems is parallel composition. Typically it generates a new transition system whose states are pairs $x||y$ of component system’s states. This suggests that the carrier of the composed system is the binary product of the carrier of its components. We thus study the case where the structure functor $T$ is ×$\times:\mathbb{B}^{2}\to\mathbb{B}$.
In this section we illustrate some liftings of the binary product functor, and show that they can be expressed by the single-parameter $2$-codensity lifting. It is easy to apply results here to the multiple-parameter codensity lifting because it is the intersection of single-parameter one.
@@ 5.1.1 Product functor on the total category
Suppose that the base category $\mathbb{B}$ has a binary product functor ×$\times:\mathbb{B}^{2}\to\mathbb{B}$. Since $p:\mathbb{E}\to\mathbb{B}$ is a ${\mathbf{CLat}_{\sqcap}}$-fibration, the functor ×˙$\mathbin{\dot{\times}}:\mathbb{E}^{2}\to\mathbb{E}$ defined by
@@
is a fibered lifting of $\times$ along $p^{2},p$. It also preserves all meets. In fact, it is the binary product functor on $\mathbb{E}$ [20, Proposition 9.2.1].
@@
Let $(\Omega,\mathbf{\Omega}):Id_{1}\to(p:\mathbb{E}\to\mathbb{B})$ be a ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values such that $\mathbb{B}$ has binary products, and $R^{p,\mathbf{\Omega}}\circ L^{p,\mathbf{\Omega}}=\mathrm{id}$. Then for any modality $\sigma:\Omega\times\Omega\to\Omega$, we have ×˙=[×$\mathbin{\dot{\times}}=[\times,\sigma]^{\mathbf{\Omega}}$ if and only if [×$[\times,\sigma]^{\mathbf{\Omega}}(\mathbf{\Omega},\mathbf{\Omega})=\mathbf{\Omega}\mathbin{\dot{\times}}\mathbf{\Omega}$. ∎
@@ 5.1.2 Metric lifting of the binary product by a modality
Here we focus on the ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values   $(\mathbb{I},(\mathbb{I},d_{\mathbb{I}})):Id_{1}\to U_{\mathbf{PMet}}$ where $\mathbb{I}$ is the interval $[0,1]$ and $(\mathbb{I},d_{\mathbb{I}})\in\mathbf{PMet}$ is the Euclidean distance over the interval on it. We will use it for quantitative bisimulations in §7.1 and §7.3.
@@
Let $\sigma\colon\mathbb{I}\times\mathbb{I}\to\mathbb{I}$ be a function. For each pseudometric spaces $(X,d_{X}),(Y,d_{Y})$, we define the function ×$\times^{\sigma}(d_{X},d_{Y})\colon(X\times Y)^{2}\to\mathbb{I}$ by
@@
Putting $f(a)\coloneqq 1-a$, the proposition above implies [×$[\times,{\sigma_{\otimes}}]^{d_{\mathbb{I}}}=[\times,{\sigma_{\oplus}}]^{d_{\mathbb{I}}}$ and [×$[\times,\sigma_{\land}]^{d_{\mathbb{I}}}=[\times,\sigma_{\lor}]^{d_{\mathbb{I}}}$ because ${\sigma_{\otimes}}=f\circ{\sigma_{\oplus}}\circ(f\times f)$ and $\sigma_{\land}=f\circ\sigma_{\lor}\circ(f\times f)$ hold,
6 Codensity Lifting of One-Step Composition
As discussed at the end of §3, our focus is on modalities that lifts one-step composition operations for coalgebras. Technically, this is equivalent to the liftability of distributive law $\lambda\colon T\circ F^{N}\to F\circ T$ with respect to codensity liftings of $F,T$.
@@ 6 Codensity Lifting of One-Step Composition
Let $((\Omega,\mathbf{\Omega}):Id_{A}\to(p:\mathbb{E}\to\mathbb{B}),F,\tau)$ be codensity bisimulation data, and $(T,\lambda)$ be an $N$-ary one-step composition on $F$-coalgebras. We say that a modality $\sigma\colon T\circ\langle\Omega\rangle_{N}\to\Omega$ lifts $\lambda$ along $p$ if $\lambda$ is liftable with respect to $[T,\sigma]^{\mathbf{\Omega}}\circ([F,\tau]^{\mathbf{\Omega}})^{N}$ and $[F,\tau]^{\mathbf{\Omega}}\circ[T,\sigma]^{\mathbf{\Omega}}$.
@@
Below we provide a sufficient condition for the modality $\sigma$ to lift a distributive law. Our approach employs the lifting of the distributive law $\lambda$ by the 2-functor $\mathrm{Sp}$ introduced in Theorem 4.5, and the decomposed definition of the generalized codensity lifting (5).
@@
This enables us to interchange the center of two codensity liftings that meets in the middle of their composition (below we put $R\triangleq R^{p,\mathbf{\Omega}},L\triangleq L^{p,\mathbf{\Omega}},R_{N}\triangleq R^{p^{N},\langle\mathbf{\Omega}\rangle_{N}},L_{N}\triangleq L^{p^{N},\langle\mathbf{\Omega}\rangle_{N}}$):
@@
Let $((\Omega,\mathbf{\Omega}):Id_{A}\to(p:\mathbb{E}\to\mathbb{B}),F,\tau)$ be codensity bisimulation data and $(T,\lambda)$ be an $N$-ary composition operation on $F$-coalgebras. Then a modality $\sigma$ lifts $\lambda$ along $p$ if the following conditions hold.
@@
We investigate the second condition. First, it can be expressed as “$\mathrm{Sp}^{A}(T,\sigma)\circ L^{p^{N},\langle\mathbf{\Omega}\rangle_{N}}(P)$ is approximating to $[F,\tau]^{\mathbf{\Omega}}$ for each $P\in\mathbb{E}^{N}$” by the concept of approximating families introduced by Komorida et al. [24] for expressivity of coalgebraic modal logics. Note that they consider a restricted setting such that $\mathbf{\Omega}$ is constant on $A$.
@@
Assume that each hom-poset $(\mathbb{B}(X,\Omega),\leq_{X})$ is a complete lattice, and we write $\bigvee$ for the join. Then $(X,S)\in\mathrm{Sp}^{A}(\mathbb{B},\Omega)$ is approximating to $[F,\tau]^{\mathbf{\Omega}}$ if the following conditions hold: for each $a\in A$,
@@
Example 6.5 (Bisimulation for ${\mathcal{P}}$-coalgebras).
This example aims for showing the composition of the standard bisimulations.
Codensity bisimulation data
We take the ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values $(2,\mathrm{Eq}_{2}):Id_{1}\to U_{\mathbf{EqRel}}$ and the covariant powerset functor ${\mathcal{P}}$ (Example 3.1) with the modality $\diamond\colon{\mathcal{P}}2\to 2$ (Example 4.2) to form codensity bisimulation data. We have $[{\mathcal{P}},\diamond]^{\mathrm{Eq}_{2}}=\dot{\mathcal{P}}$ (Example 3.3).
Binary one-step composition operation for ${\mathcal{P}}$-coalgebras
Modality for $\times$
We adopt the modality :2×22$\land$:2$\times$ 2$\to$ 2∧ : 2 × 2 → 2 in Example 5.2. The codensity lifting [×$[\times,\land]^{\mathrm{Eq}_{2}}$ coincides with the binary product of predicates as in Example 5.2.
Proposition 6.6.
This proposition, together with Corollary 3.10, implies that for two ${\mathcal{P}}$-coalgebras $c_{1}:X_{1}\to{\mathcal{P}}(X_{1})$ and $c_{2}:X_{2}\to{\mathcal{P}}(X_{2})$, if $x_{i}$ and $x_{i}^{\prime}$ are $[{\mathcal{P}},\diamond]^{\mathrm{Eq}_{2}}$-bisimilar in $c_{i}$ ($i\in\{1,2\}$), the pairs $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ are $[{\mathcal{P}},\diamond]^{\mathrm{Eq}_{2}}$-bisimilar in the composed ${\mathcal{P}}$-coalgebra $\lambda^{{\mathcal{P}}}\circ(c_{1}\times c_{2})$.
The proof of Proposition 6.6 is the following: $\wedge$ satisfies two conditions in Theorem 6.2: 1) For each $A,B\subseteq 2$, $\diamond A\land\diamond B=\diamond\{a\land b\mid a\in A,b\in B\}$. 2) As discussed before Definition 6.3, it’s equivalent to approximating objects $\{\land\circ(k_{P}\times k_{Q})\mid k_{P}\colon P\to\mathrm{Eq}_{2},k_{Q}\colon Q\to\mathrm{Eq}_{2}\}$ for each $P,Q\in\mathbf{EqRel}$. We can prove it by Proposition 6.4 with the ordered object $(2,\leq)$ where $\leq$ is the order satisfying $\mathrm{false}\leq\mathrm{true}$. One easily verifies the first and second condition in Proposition 6.4. For the third condition, for each $a\in A$ and $k\colon[\times,\land]^{\mathrm{Eq}_{2}}(P,Q)\to\mathrm{Eq}_{2}$, we define $S_{k}^{\prime}\coloneqq\{\land\circ(\chi_{[x]_{P}}\times\chi_{[y]_{Q}})\mid k(x,y)=1\}$ where $\chi_{[x]_{P}},\chi_{[y]_{Q}}$ are characteristic functions. Then $pk=\bigvee_{k^{\prime}\in S^{\prime}_{k}}k^{\prime}$ holds.
Example 6.7.
Codensity bisimulation data
They are set-up in Example 4.3. We package the ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values $(\Omega,\mathbf{\Omega}):Id_{\Sigma\uplus\{\epsilon\}}\to U_{\mathbf{EqRel}}$, the behavior functor ${F_{\mathrm{da}}}$ and the modality $\tau$ there into codensity bisimulation data.
Binary one-step composition operation for ${F_{\mathrm{da}}}$-coalgebras
We adopt the one (×$(\times,\lambda^{\mathrm{da}})$ in Example 3.6.
Modality for $\times$
We adopt the modality $\sigma$ given by $\sigma_{a}\triangleq\land$ (the logical conjunction; here $a\in\Sigma\uplus\{\epsilon\}$). The binary codensity lifting [×$[\times,\land]^{\mathrm{Eq}_{2}}$ is equal to that in Example 5.2, so that [×$[\times,\land]^{\mathrm{Eq}_{2}}=\mathbin{\dot{\times}}$.
Proposition 6.8.
As a consequence, given deterministic automata $c_{1}:X_{1}\to{F_{\mathrm{da}}}(X_{1})$ and $c_{2}:X_{2}\to{F_{\mathrm{da}}}(X_{2})$, if $x_{i},x_{i}^{\prime}\in X_{i}$ are language equivalent in $c_{i}$ ($i\in\{1,2\}$), the pairs $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ are also language equivalent in the product automaton $\lambda^{\mathrm{da}}\circ(c_{1}\times c_{2})$.
The proof of Proposition 6.8 is as follows. The modality $\land$ satisfies two conditions in Theorem 6.2: 1) For each $(t,\rho),(t^{\prime},\rho^{\prime})\in 2\times 2^{\Sigma}$, both $\land\circ\tau_{a}$ and $\tau_{\epsilon}\circ(2\times\land^{\Sigma})\circ\lambda^{\mathrm{da}}$ map $((t,\rho),(t^{\prime},\rho^{\prime}))$ to $t\land t$ if $a=\epsilon$ and $\rho(a)\land\rho^{\prime}(a)$ otherwise. 2) We can prove it by Proposition 6.4. The second and third conditions are the same as Example. 6.5. The first condition in Proposition 6.4 holds because for each $S^{\prime}\subseteq\mathbf{Set}(X^{\prime},2)$ and $(t,\rho)\in 2\times X^{\prime\Sigma}$, both $\tau_{a}\circ 2\times(\bigvee_{f\in S^{\prime}}f)^{\Sigma}$ and $\bigvee_{f\in S^{\prime}}\tau_{a}\circ 2\times f^{\Sigma}$ map $(t,\rho)$ to $t$ if $a=\epsilon$ and $\bigvee_{f\in S^{\prime}}f(\rho(a))$ otherwise.

### v2

@@ 5 Codensity Liftings of structure Functors
We use the generalized codensity lifting to lift the structure functor. Let $(\Omega,\mathbf{\Omega})\colon Id_{A}\to(p\colon\mathbb{E}\to\mathbb{B})$ be a ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values. Its $N$-fold tupling $(\langle\Omega\rangle_{N},\langle\mathbf{\Omega}\rangle_{N})\colon Id_{A}\to(p^{N}\colon\mathbb{E}^{N}\to\mathbb{B}^{N})$ is again a ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values; here $\langle\Omega\rangle_{N}$ denotes the tupling $\langle\Omega,\cdots,\Omega\rangle\colon A\to\mathbb{B}^{N}$ of $N$-fold copies of $\Omega$ (and the same for $\langle\mathbf{\Omega}\rangle_{N}$). We call the codensity lifting of a structure functor $T\colon\mathbb{B}^{N}\to\mathbb{B}$ with $\sigma\colon T\circ\langle\Omega\rangle_{N}\to\Omega$ along $p^{N},p$ the $N$-codensity lifting. Concretely,
@@ 5.1 Product Functors by Codensity Liftings
One of the most fundamental operators for composing processes and transition systems is parallel composition. Typically it generates a new transition system whose states are pairs $x||y$ of component system’s states. This suggests that the carrier of the composed system is the binary product of the carrier of its components. We thus study the case where the structure functor $T$ is ×$\times\colon\mathbb{B}^{2}\to\mathbb{B}$.
As a side note, we can also handle other structure functors including coproducts, thanks to the generality of §4. Moreover, our framework facilitates the investigation of various modalities, even when we fix $T$ to the product functor, as illustrated in §5.1.2. This enables us to accommodate various non-trivial liftings for achieving compositionality, particulary in quantitative analysis (§7).
In this section we illustrate some liftings of the binary product functor, and show that they can be expressed by the single-parameter $2$-codensity lifting. It is easy to apply results here to the multiple-parameter codensity lifting because it is the intersection of the single-parameter one.
@@ 5.1.1 Product functor on the total category
Suppose that the base category $\mathbb{B}$ has a binary product functor ×$\times\colon\mathbb{B}^{2}\to\mathbb{B}$. Since $p\colon\mathbb{E}\to\mathbb{B}$ is a ${\mathbf{CLat}_{\sqcap}}$-fibration, the functor ×˙$\mathbin{\dot{\times}}\colon\mathbb{E}^{2}\to\mathbb{E}$ defined by
@@
is a fibered lifting of $\times$ along $p^{2},p$. It also preserves all meets. In fact, it is the binary product functor on $\mathbb{E}$ [19, Proposition 9.2.1].
@@
Let $(\Omega,\mathbf{\Omega})\colon Id_{1}\to(p\colon\mathbb{E}\to\mathbb{B})$ be a ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values such that $\mathbb{B}$ has binary products, and $R^{p,\mathbf{\Omega}}\circ L^{p,\mathbf{\Omega}}=\mathrm{id}$. Then for any modality $\sigma\colon\Omega\times\Omega\to\Omega$, we have ×˙=[×$\mathbin{\dot{\times}}=[\times,\sigma]^{\mathbf{\Omega}}$ if and only if [×$[\times,\sigma]^{\mathbf{\Omega}}(\mathbf{\Omega},\mathbf{\Omega})=\mathbf{\Omega}\mathbin{\dot{\times}}\mathbf{\Omega}$. ∎
@@ 5.1.2 Metric lifting of the binary product by a modality
Here we focus on the ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values   $(\mathbb{I},(\mathbb{I},d_{\mathbb{I}}))\colon Id_{1}\to U_{\mathbf{PMet}}$ where $\mathbb{I}$ is the interval $[0,1]$ and $(\mathbb{I},d_{\mathbb{I}})\in\mathbf{PMet}$ is the Euclidean distance over the interval on it. We will use it for quantitative bisimulations in §7.1 and §7.3.
@@
Let $\sigma\colon\mathbb{I}\times\mathbb{I}\to\mathbb{I}$ be a function. Given pseudometric spaces $(X,d_{X}),(Y,d_{Y})$, we define the function ×$\times^{\sigma}(d_{X},d_{Y})\colon(X\times Y)^{2}\to\mathbb{I}$ by
@@
Putting $f(a)\coloneqq 1-a$, the proposition above implies [×$[\times,{\sigma_{\otimes}}]^{d_{\mathbb{I}}}=[\times,{\sigma_{\oplus}}]^{d_{\mathbb{I}}}$ and [×$[\times,\sigma_{\land}]^{d_{\mathbb{I}}}=[\times,\sigma_{\lor}]^{d_{\mathbb{I}}}$ because ${\sigma_{\otimes}}=f\circ{\sigma_{\oplus}}\circ(f\times f)$ and $\sigma_{\land}=f\circ\sigma_{\lor}\circ(f\times f)$ hold.
6 Codensity Lifting of One-Step Composition Operations
As discussed at the end of §3, our focus is on modalities that lift one-step composition operations for coalgebras. Technically, this is equivalent to the liftability of distributive law $\lambda\colon T\circ F^{N}\to F\circ T$ with respect to codensity liftings of $F,T$.
@@ 6 Codensity Lifting of One-Step Composition Operations
Let $((\Omega,\mathbf{\Omega})\colon Id_{A}\to(p\colon\mathbb{E}\to\mathbb{B}),F,\tau)$ be codensity bisimulation data, and $(T,\lambda)$ be an $N$-ary one-step composition on $F$-coalgebras, see §3.2 for the definition. We say that a modality $\sigma\colon T\circ\langle\Omega\rangle_{N}\to\Omega$ lifts $\lambda$ along $p$ if $\lambda$ is liftable with respect to $[T,\sigma]^{\mathbf{\Omega}}\circ([F,\tau]^{\mathbf{\Omega}})^{N}$ and $[F,\tau]^{\mathbf{\Omega}}\circ[T,\sigma]^{\mathbf{\Omega}}$.
@@
6.1 Sufficient Condition for Liftability
We provide a sufficient condition for the modality $\sigma$ to lift a distributive law. Our approach employs the lifting of the distributive law $\lambda$ by the 2-functor $\mathrm{Sp}$ introduced in Theorem 4.5, and the decomposed definition of the generalized codensity lifting (5).
@@ 6.1 Sufficient Condition for Liftability
This lifting enables us to interchange the center of two codensity liftings that meets in the middle of their composition (below we put $R\coloneq R^{p,\mathbf{\Omega}},L\coloneq L^{p,\mathbf{\Omega}},R_{N}\coloneq R^{p^{N},\langle\mathbf{\Omega}\rangle_{N}},L_{N}\coloneq L^{p^{N},\langle\mathbf{\Omega}\rangle_{N}}$):
@@
Let $((\Omega,\mathbf{\Omega})\colon Id_{A}\to(p\colon\mathbb{E}\to\mathbb{B}),F,\tau)$ be codensity bisimulation data and $(T,\lambda)$ be an $N$-ary composition operation on $F$-coalgebras. Then a modality $\sigma$ lifts $\lambda$ along $p$ if the following conditions hold.
@@
We investigate the second condition. First, it can be expressed as “$\mathrm{Sp}^{A}(T,\sigma)\circ L^{p^{N},\langle\mathbf{\Omega}\rangle_{N}}(P)$ is approximating to $[F,\tau]^{\mathbf{\Omega}}$ for each $P\in\mathbb{E}^{N}$” by the concept of approximating families introduced by Komorida et al. [24] for expressivity of coalgebraic modal logics (in a restricted setting where $\mathbf{\Omega}$ is constant on $A$).
@@
Assume that each hom-poset $(\mathbb{B}(X,\Omega),\leq_{X})$ is a complete lattice, and write $\bigvee$ for the join. Then $(X,S)\in\mathrm{Sp}^{A}(\mathbb{B},\Omega)$ is approximating to $[F,\tau]^{\mathbf{\Omega}}$ if all of the following hold, for each $a\in A$:
@@
6.1.1 Example: Bisimulation for ${\mathcal{P}}$-coalgebras
This example aims to show the composition of standard bisimulations.
Codensity bisimulation data
We take the ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values $(2,\mathrm{Eq}_{2})\colon Id_{1}\to U_{\mathbf{EqRel}}$ and the covariant powerset functor ${\mathcal{P}}$ (Example 3.1) with the modality $\diamond\colon{\mathcal{P}}2\to 2$ (Example 4.2) to form codensity bisimulation data. We have $[{\mathcal{P}},\diamond]^{\mathrm{Eq}_{2}}=\dot{\mathcal{P}}$ (Example 3.3).
Binary one-step composition operation for ${\mathcal{P}}$-coalgebras
Modality for $\times$
[…]
Proposition 6.7.
This proposition, together with Corollary 3.10, implies that for two ${\mathcal{P}}$-coalgebras $c_{1}\colon X_{1}\to{\mathcal{P}}(X_{1})$ and $c_{2}\colon X_{2}\to{\mathcal{P}}(X_{2})$, if $x_{i}$ and $x_{i}^{\prime}$ are $[{\mathcal{P}},\diamond]^{\mathrm{Eq}_{2}}$-bisimilar in $c_{i}$ ($i\in\{1,2\}$), the pairs $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ are $[{\mathcal{P}},\diamond]^{\mathrm{Eq}_{2}}$-bisimilar in the composed ${\mathcal{P}}$-coalgebra $\lambda^{{\mathcal{P}}}\circ(c_{1}\times c_{2})$.
The proof of Proposition 6.7 is the following: $\wedge$ satisfies two conditions in Theorem 6.2: 1) For each $A,B\subseteq 2$, $\diamond A\land\diamond B=\diamond\{a\land b\mid a\in A,b\in B\}$. 2) As discussed before Definition 6.3, it is equivalent to approximating objects $\{\land\circ(k_{P}\times k_{Q})\mid k_{P}\colon P\to\mathrm{Eq}_{2},k_{Q}\colon Q\to\mathrm{Eq}_{2}\}$ for each $P,Q\in\mathbf{EqRel}$. We can prove it by Proposition 6.4 with the ordered object $(2,\leq)$ where $\leq$ is the order satisfying $\mathrm{false}\leq\mathrm{true}$. One easily verifies the first and second condition in Proposition 6.4. For the third condition, for each $a\in A$ and $k\colon[\times,\land]^{\mathrm{Eq}_{2}}(P,Q)\to\mathrm{Eq}_{2}$, we define $S_{k}^{\prime}\coloneqq\{\land\circ(\chi_{[x]_{P}}\times\chi_{[y]_{Q}})\mid k(x,y)=1\}$ where $\chi_{[x]_{P}},\chi_{[y]_{Q}}$ are characteristic functions. Then $pk=\bigvee_{k^{\prime}\in S^{\prime}_{k}}k^{\prime}$ holds.
6.1.2 Example: Language Equivalence of Deterministic Automata
Codensity bisimulation data
They are set-up in Example 4.3. We package the ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values $(\Omega,\mathbf{\Omega})\colon Id_{\Sigma\uplus\{\epsilon\}}\to U_{\mathbf{EqRel}}$, the behavior functor ${F_{\mathrm{da}}}$ and the modality $\tau$ there into codensity bisimulation data.
Binary one-step composition operation for ${F_{\mathrm{da}}}$-coalgebras
We adopt the one (×$(\times,\lambda^{\mathrm{da}})$ in Example 3.6.
Modality for $\times$
We adopt the modality $\sigma$ given by $\sigma_{a}\coloneq\land$ :- ∧ (the logical conjunction; here $a\in\Sigma\uplus\{\epsilon\}$). The binary codensity lifting [×$[\times,\land]^{\mathrm{Eq}_{2}}$ is equal to that in Example 5.2, so that [×$[\times,\land]^{\mathrm{Eq}_{2}}=\mathbin{\dot{\times}}$.
Proposition 6.8.
As a consequence, given deterministic automata $c_{1}\colon X_{1}\to{F_{\mathrm{da}}}(X_{1})$ and $c_{2}\colon X_{2}\to{F_{\mathrm{da}}}(X_{2})$, if $x_{i},x_{i}^{\prime}\in X_{i}$ are language equivalent in $c_{i}$ ($i\in\{1,2\}$), the pairs $(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2}^{\prime})$ are also language equivalent in the product automaton $\lambda^{\mathrm{da}}\circ(c_{1}\times c_{2})$.
The proof of Proposition 6.8 is as follows. The modality $\land$ satisfies two conditions in Theorem 6.2: 1) For each $(t,\rho),(t^{\prime},\rho^{\prime})\in 2\times 2^{\Sigma}$, both $\land\circ\tau_{a}$ and $\tau_{\epsilon}\circ(2\times\land^{\Sigma})\circ\lambda^{\mathrm{da}}$ map $((t,\rho),(t^{\prime},\rho^{\prime}))$ to $t\land t$ if $a=\epsilon$ and $\rho(a)\land\rho^{\prime}(a)$ otherwise. 2) We can prove it by Proposition 6.4. The second and third conditions are the same as §6.1.1. The first condition in Proposition 6.4 holds because for each $S^{\prime}\subseteq\mathbf{Set}(X^{\prime},2)$, both $\tau_{a}\circ 2\times(\bigvee_{f\in S^{\prime}}f)^{\Sigma}$ and $\bigvee_{f\in S^{\prime}}\tau_{a}\circ 2\times f^{\Sigma}$ map $(t,\rho)\in 2\times X^{\prime\Sigma}$ to $t$ if $a=\epsilon$, and $\bigvee_{f\in S^{\prime}}f(\rho(a))$ otherwise.